# 2cm2cm1cm1cm

### v1

@@ 4.2.1 Essential spectrum
Observing that $A_{1,2}(k)$, $A_{2,1}(k)$, $A_{2,2}(k)$ and $P_{x}$ are all rank-one linear transformations, we can apply the stability of the essential spectrum under compact perturbations (see [49, Corollary 8.16]) to conclude that
@@ 4.2.2 Discrete spectrum
for all $k\in\mathbb{T}^{2}$, with equality only at $k=0$. See Equation (27) and Corollary 4.2. By direct computations101010(63) is clearly true for $h_{b}=0$. Take $h_{b}>0$. Using $\cos\left(\theta\right)=2\cos^{2}\left(\theta/2\right)-1$ and (30), one verifies that (63) is equivalent to $h_{b}\left(4-2x^{2}-2y^{2}\right)\leq 4-2\left(x+y\right)$ for $x,y\in\left[0,1\right]$. Since $\inf_{x\in\left[0,1\right]}\left\{h_{b}x^{2}-x\right\}=-1/(4h_{b})$ for $h_{b}\geq 1/2$ and $\inf_{x\in\left[0,1\right]}\left\{h_{b}x^{2}-x\right\}=h_{b}-1$ for $h_{b}\in(0,1/2]$, we deduce that (63) holds true iff $h_{b}\in[0,1/2]$., one verifies that this amounts to take $h_{b}$ in the interval $[0,1/2]$. This means that we consider a regime where the boson mass is at least the mass of the two fermions, as physically expected for cuprate superconductors, see [21, Section 3.1].
@@
(see (28) and (30)) defines a positive and continuous function that vanishes at $p=0$ only, one deduces from (70) that $\varphi=0$ also when $k=0$. Therefore, for any fixed $k\in\mathbb{T}^{2}$, $\varphi=0$ and so, (68) combined with (38) yields in this case
$A_{1,2}(k)z\doteq\hat{\upsilon}(k)\mathfrak{d}(k)z=0\text{ },$
i.e., $z=0$. Thus, we arrive at $(\varphi,z)=(0,0)$, which contradicts the fact that $(\varphi,z)$ is a non-zero vector. Therefore, if $\mathfrak{b}(k)$ is an eigenvalue of $A(k)$ then we must have $\hat{\upsilon}(k)=0$.
@@ 4.2.3 Bottom of the spectrum
Proof. Assume that $k\neq 0$. Recall that $\mathfrak{z}(k)$ is defined in Corollary 4.2. From Corollaries 4.2 and 4.6, the interval $(-\infty,\mathfrak{z}(k))$ contains at most one eigenvalue of $A(k)$. By Corollary 4.5, the eigenvalue is non-degenerate, if it exists. If $\hat{\upsilon}(k)=0$, we know from Proposition 4.3 (ii) that $\mathfrak{b}(k)$ is such an eigenvalue. Recall that for $k\neq 0$ one has that $\mathfrak{b}(k)<\mathfrak{z}(k)$, because $h_{b}\in[0,1/2]$. Now, suppose that $\hat{\upsilon}(k)\neq 0$. When $\mathfrak{b}(k)\leq x<\mathfrak{z}(k)$, we have
@@
Take again some $k_{0}\in\mathbb{S}^{2}$. Assume that $\hat{\upsilon}$ is of class $C^{d}$ on $\mathbb{S}^{2}\subseteq\mathbb{R}^{2}$ with $d\in\mathbb{N}\cup\{\omega,a\}$. Let $\vartheta=(k_{0},\mathrm{E}(k_{0}))\in\mathcal{O}$. Using Theorems 4.4 and 4.7 as well as Equation (77) and the (operator norm) continuity of the mapping
@@
(in operator norm). Hence, $\lambda\in\sigma(A(k_{0}))$, for otherwise $A(k_{n})-\lambda_{n}\mathfrak{1}$ would be invertible [^12] Recall that the set of invertible operators on a Banach space $X$ is an open subset of $\mathcal{B}(X)$, with respect to the operator norm. See, e.g., [44, Theorem 1.4]. for sufficiently large $n$. Thus, $\mathcal{K}$ is a closed set and, as a consequence, for any $s\notin\mathcal{K}$, there is $\varepsilon>0$ such that
@@ 4.3.1 The characteristic equation in the hard-core limit
In the following it is technically convenient to assume that $\mathfrak{d}(k)\notin\mathbb{C}\mathfrak{s}$ for all $k\in\mathbb{T}^{2}$. Notice that this holds true iff $\mathfrak{p}\notin\mathbb{C}\mathfrak{e}_{0}$, i.e., $r_{\mathfrak{p}}>0$. See discussions around Equation (13).
@@
with $R_{\mathfrak{d},\mathfrak{d}}R_{\mathfrak{s},\mathfrak{s}}-\left|R_{\mathfrak{s},\mathfrak{d}}\right|^{2}\geq 0$. Moreover, if $r_{\mathfrak{p}}>0$ (see (13)), then the inequality is strict.
@@ 4.3.2 Hard-core dispersion relation of bound pairs of lowest energy
We give in the next theorem a hard-core limit version of Theorems 4.4, 4.7 and 4.8. To this end, recall that $\mathrm{E}\left(k\right)$, $k\in\mathbb{T}^{2}$, are given by Theorem 4.7 as a family of non-degenerate eigenvalues. This family depends upon the parameter $\mathrm{U}\in\mathbb{R}_{0}^{+}$ and we thus use here the notation $\mathrm{E}(\mathrm{U},k)\equiv\mathrm{E}(k)$. This defines a function $\mathrm{E}:\mathbb{R}_{0}^{+}\times\mathbb{T}^{2}\rightarrow\mathbb{R}$.
@@
Finally, assume that $r_{\mathfrak{p}}>0$, that is, $\mathfrak{p}\notin\mathbb{C}\mathfrak{e}_{0}$. For $k\neq 0$, we deduce from (99) and Lemma 4.11 that
@@
keeping in mind that $\mathfrak{b}\left(0\right)=\mathfrak{z}(0)=0$. Then, we infer from Theorems 4.4 and 4.7 that the following equality must be true:
@@
Under the additional conditions that $r_{\mathfrak{p}}>0$ (i.e., $\mathfrak{p}\notin\mathbb{C}\mathfrak{e}_{0}$) and $\hat{\upsilon}(0)\neq 0$, Assertion (iv) remains valid with $(-\pi,\pi)^{2}$ instead of $\mathbb{S}^{2}$, i.e., by including the zero quasi-momentum case $k=0$. Also, Assertion (iii) holds true for $k=0$. In fact, under these further conditions, we know from Assertion (v) that $\mathrm{E}(\infty,0)<\mathfrak{z}(0)=0$.
@@
By combining Equation (111) with the mean value theorem, we see that $\mathfrak{F}$ is Lipschitz equicontinuous. In particular this family of functions is equicontinuous. Moreover, it follows from (110) that $\mathfrak{F}$ is bounded in the supremum norm. We can hence apply the (Arzelà-) Ascoli theorem [38, Theorem A5], according to which $\partial_{x}\Phi(\mathrm{U}_{n},k_{0},\cdot)$, when restricted to the compact interval $[x_{0}-\delta,x_{0}+\delta]$, converges uniformly along some subsequence. Assume, for simplicity and without loss of generality, that the (full) sequence of functions converges itself. By Equation (102), recall that $\Phi(\mathrm{U},\cdot,\cdot)$ converges pointwise to the function $\Phi(\infty,\cdot,\cdot)$ on $\mathcal{O}$, which is is the real-valued function defined by (101). Then, by [47, Theorem 7.17],
@@
for every $(k,x)\in\mathcal{O}$. In the same way, we invoke Equations (108) and (109) together with the mean value theorem and the (Arzelà-) Ascoli theorem [38, Theorem A5] to deduce that
@@
This statement remains valid for $k=0$ provided that $\hat{\upsilon}(0)\neq 0$ and $r_{\mathfrak{p}}>0$ (i.e., $\mathfrak{p}\notin\mathbb{C}\mathfrak{e}_{0}$).
@@
Note that $\mathrm{E}(\infty,k)\leq\mathfrak{b}\left(k\right)<\mathfrak{z}\left(k\right)$ for any $k\neq 0$. (When $\hat{\upsilon}(0)\neq 0$ and $\mathfrak{p}\notin\mathbb{C}\mathfrak{e}_{0}$, we also have that $\mathrm{E}(\infty,0)<\mathfrak{b}(0)=\mathfrak{z}(0)$.) On the other hand, by Proposition 5.14, $\{(A_{1,1}(\mathrm{U},k)-\mathrm{E}(\infty,k)\mathfrak{1})^{-1}\}_{\mathrm{U}\geq 0}$ is a decreasing family of positive operators and, by Proposition 5.15, it converges strongly as $\mathrm{U}\rightarrow\infty$. Consequently, $(A_{1,1}(\mathrm{U},k)-\mathrm{E}(\mathrm{U},k)\mathfrak{1})^{-1}$ also converges strongly.
@@ 4.4 Spectral Gap and Anderson Localization
for any Hubbard coupling constant $\mathrm{U}\in\mathbb{R}_{0}^{+}$. When $r_{\mathfrak{p}}>0$ (i.e., $\mathfrak{p}\notin\mathbb{C}\mathfrak{e}_{0}$) and $\hat{\upsilon}(0)\neq 0$, this quantity turns out to be strictly positive, uniformly with respect to the parameter $\mathrm{U}$:
@@
Fix $h_{b}\in[0,1/2]$. If $r_{\mathfrak{p}}>0$ (i.e., $\mathfrak{p}\notin\mathbb{C}\mathfrak{e}_{0}$) and $\hat{\upsilon}(0)\neq 0$ then
@@
$\mathfrak{d}(k)(p)\doteq\mathfrak{\hat{p}}(k+p)\ ,\qquad p\in\mathbb{T}^{2}\ ,$ (120)
@@
Proof. By (26) and (120), for any $p,k\in\mathbb{T}^{2}$, note that
$\mathfrak{d}(k)(p)={\sum_{x\in\mathbb{Z}^{2}}}\,\mathrm{e}^{i(k+p)\cdot x}\mathfrak{p}(x)=\mathcal{F}[\mathrm{e}^{ik\cdot(x)}\mathfrak{p}(x)](p)\ ,$
where $\mathrm{e}^{ik\cdot(x)}\mathfrak{p}(x)$ stands for the function $x\mapsto\mathrm{e}^{ik\cdot(x)}\mathfrak{p}(x)$. It follows that
$\mathcal{F}^{-1}\left[(A_{1,1}(\mathrm{U},k)-\mathrm{E}(\mathrm{U},k)\mathfrak{1})^{-1}\mathfrak{d}(k)\right]$ $=(\mathcal{F}^{-1}A_{1,1}(\mathrm{U},k)\mathcal{F}-\mathrm{E}(\mathrm{U},k)\mathfrak{1})^{-1}\mathcal{F}^{-1}\mathfrak{d}(k)$
$={\sum_{y\in\mathbb{Z}^{2}}}\,\mathrm{e}^{ik\cdot y}\mathfrak{p}(y)(\mathcal{F}^{-1}A_{1,1}(\mathrm{U},k)\mathcal{F}-\mathrm{E}(\mathrm{U},k)\mathfrak{1})^{-1}\mathfrak{e}_{y}\ .$ (122)
@@
$\left|\psi_{\mathrm{U},k}\left(x\right)\right|$ $=$ $\left|\left\langle\mathfrak{e}_{x},\psi_{\mathrm{U},k}\right\rangle\right|\leq\left|\hat{\upsilon}\left(k\right)\right|{\sum_{y\in\mathbb{Z}^{2}}}\left|\mathfrak{p}\left(y\right)\right|\left|\left\langle\mathfrak{e}_{x},\left(\mathcal{F}^{-1}A_{1,1}\left(\mathrm{U},k\right)\mathcal{F}-\mathrm{E}(\mathrm{U},k)\mathfrak{1}\right)^{-1}\mathfrak{e}_{y}\right\rangle\right|$ (127)
$\leq$ $\frac{\left|\hat{\upsilon}\left(k\right)\right|}{\inf_{\mathrm{U}\in\mathbb{R}_{0}^{+}}\min_{k\in\mathbb{T}^{2}}\mathfrak{g}\left(\mathrm{U},k\right)-4\epsilon(\mathrm{e}^{\alpha}-1)}{\sum_{y\in\mathbb{Z}^{2}}}\left|\mathfrak{p}\left(y\right)\right|\mathrm{e}^{-\alpha\left|x-y\right|}$
$\leq$ $\frac{\left|\hat{\upsilon}\left(k\right)\right|\mathrm{e}^{-\alpha\left|x\right|}}{\inf_{\mathrm{U}\in\mathbb{R}_{0}^{+}}\min_{k\in\mathbb{T}^{2}}\mathfrak{g}\left(\mathrm{U},k\right)-4\epsilon(\mathrm{e}^{\alpha}-1)}{\sum_{y\in\mathbb{Z}^{2}}}\left|\mathfrak{p}\left(y\right)\right|\mathrm{e}^{\alpha\left|y\right|}$

### v2

@@ 4.3.1 Essential spectrum
Observing that $A_{1,2}(k)$, $A_{2,1}(k)$, $A_{2,2}(k)$ and $P_{x}$ are all rank-one linear transformations, we can apply the stability of the essential spectrum under compact perturbations (see [95, Corollary 8.16]) to conclude that
@@ 4.3.2 Discrete spectrum
for all $k\in\mathbb{T}^{2}$, with equality only at $k=0$. See Equation (32) and Corollary 4.3. By direct computations171717(79) is clearly true for $h_{b}=0$. Take $h_{b}>0$. Using $\cos\left(\theta\right)=2\cos^{2}\left(\theta/2\right)-1$ and (35), one verifies that (79) is equivalent to $h_{b}\left(4-2x^{2}-2y^{2}\right)\leq 4-2\left(x+y\right)$ for $x,y\in\left[0,1\right]$. Since $\inf_{x\in\left[0,1\right]}\left\{h_{b}x^{2}-x\right\}=-1/(4h_{b})$ for $h_{b}\geq 1/2$ and $\inf_{x\in\left[0,1\right]}\left\{h_{b}x^{2}-x\right\}=h_{b}-1$ for $h_{b}\in(0,1/2]$, we deduce that (79) holds true iff $h_{b}\in[0,1/2]$., one verifies that this amounts to take $h_{b}$ in the interval $[0,1/2]$. This means that we consider a regime where the boson mass is at least the mass of the two fermions, as physically expected for cuprate superconductors, see [21, Section 3.1].
@@
(see (33) and (35)) defines a positive and continuous function that vanishes at $p=0$ only, one deduces from (86) that $\varphi=0$ also when $k=0$. In any case, $\varphi=0$ and so, (84) combined with (44) yields
$A_{1,2}(k)z\doteq\hat{\upsilon}(k)\mathfrak{d}(k)z=0\text{ }.$
Since $\mathfrak{d}(k)\neq 0$ and $\hat{\upsilon}(k)\neq 0$, we must have that $z=0$.Thus, we arrive at $(\varphi,z)=(0,0)$, which contradicts the fact that $(\varphi,z)$ is a non-zero vector. Therefore, if $\mathfrak{b}(k)$ is an eigenvalue of $A(k)$ then we must have $\hat{\upsilon}(k)=0$.
@@ 4.3.3 Bottom of the spectrum
Proof. Assume that $k\neq 0$. Recall that $\mathfrak{z}(k)$ is defined in Corollary 4.3. From Corollaries 4.3 and 4.7, the interval $(-\infty,\mathfrak{z}(k))$ contains at most one eigenvalue of $A(k)$. By Corollary 4.6, the eigenvalue is non-degenerate, if it exists. If $\hat{\upsilon}(k)=0$, we know from Proposition 4.4 (ii) that $\mathfrak{b}(k)$ is such an eigenvalue. Recall that for $k\neq 0$ one has that $\mathfrak{b}(k)<\mathfrak{z}(k)$, because $h_{b}\in[0,1/2]$. Now, suppose that $\hat{\upsilon}(k)\neq 0$. When $\mathfrak{b}(k)\leq x<\mathfrak{z}(k)$, we have
@@
Take again some $k_{0}\in\mathbb{S}^{2}$. Assume that $\hat{\upsilon}$ is of class $C^{d}$ on $\mathbb{S}^{2}\subseteq\mathbb{R}^{2}$ with $d\in\mathbb{N}\cup\{\omega,a\}$. Let $\vartheta=(k_{0},\mathrm{E}(k_{0}))\in\mathcal{O}$. Using Theorems 4.5 and 4.8 as well as Equation (93) and the (operator norm) continuity of the mapping
@@
(in operator norm). Hence, $\lambda\in\sigma(A(k_{0}))$, for otherwise $A(k_{n})-\lambda_{n}\mathfrak{1}$ would be invertible [^19] Recall that the set of invertible operators on a Banach space $X$ is an open subset of $\mathcal{B}(X)$, with respect to the operator norm. See, e.g., [50, Theorem 1.4]. for sufficiently large $n$. Thus, $\mathcal{K}$ is a closed set and, as a consequence, for any $s\notin\mathcal{K}$, there is $\varepsilon>0$ such that
@@ 4.4.1 The characteristic equation in the hard-core limit
In the following it is technically convenient to assume that $\mathfrak{d}(k)\notin\mathbb{C}\mathfrak{s}$ for all $k\in\mathbb{T}^{2}$. Notice that this holds true iff $\mathfrak{p}_{1}\notin\mathbb{C}\mathfrak{e}_{0}$ or $\mathfrak{p}_{2}\notin\mathbb{C}\mathfrak{e}_{0}$, i.e., $r_{\mathfrak{p}}>0$. Indeed, recall (36), that is,
$\mathfrak{d}\left(k\right)=\mathcal{F}\left[\mathrm{e}^{ik\cdot x}\mathfrak{p}_{1}\left(x\right)+\mathrm{e}^{i\frac{k}{2}\cdot x}\mathfrak{p}_{2}\left(x\right)\right]\ ,$
where $\mathrm{e}^{ik\cdot x}\mathfrak{p}_{\sharp}(x)$ stands for the function $x\mapsto\mathrm{e}^{ik\cdot x}\mathfrak{p}_{\sharp}(x)$ with $\sharp\in\{1,2\}$. See also discussions around Equations (13)–(14).
@@
with $R_{\mathfrak{d},\mathfrak{d}}R_{\mathfrak{s},\mathfrak{s}}-\left|R_{\mathfrak{s},\mathfrak{d}}\right|^{2}\geq 0$. Moreover, if $r_{\mathfrak{p}}>0$ (see (14)–(15)), then the inequality is strict.
@@ 4.4.2 Hard-core dispersion relation of bound pairs of lowest energy
We give in the next theorem a hard-core limit version of Theorems 4.5, 4.8 and 4.9. To this end, recall that $\mathrm{E}\left(k\right)$, $k\in\mathbb{T}^{2}$, are given by Theorem 4.8 as a family of non-degenerate eigenvalues. This family depends upon the parameter $\mathrm{U}\in\mathbb{R}_{0}^{+}$ and we thus use here the notation $\mathrm{E}(\mathrm{U},k)\equiv\mathrm{E}(k)$. This defines a function $\mathrm{E}:\mathbb{R}_{0}^{+}\times\mathbb{T}^{2}\rightarrow\mathbb{R}$.
@@
Finally, assume that $r_{\mathfrak{p}}>0$, that is, $\mathfrak{p}_{1}\notin\mathbb{C}\mathfrak{e}_{0}$ or $\mathfrak{p}_{2}\notin\mathbb{C}\mathfrak{e}_{0}$.For $k\neq 0$, we deduce from (115) and Lemma 4.12 that
@@
keeping in mind that $\mathfrak{b}\left(0\right)=\mathfrak{z}(0)=0$. Then, we infer from Theorems 4.5 and 4.8 that the following equality must be true:
@@
Under the additional conditions that $r_{\mathfrak{p}}>0$ (i.e., $\mathfrak{p}_{1}\notin\mathbb{C}\mathfrak{e}_{0}$ or $\mathfrak{p}_{2}\notin\mathbb{C}\mathfrak{e}_{0}$) and $\hat{\upsilon}(0)\neq 0$, Assertion (iv) remains valid with $(-\pi,\pi)^{2}$ instead of $\mathbb{S}^{2}$, i.e., by including the zero quasi-momentum case $k=0$. Also, Assertion (iii) holds true for $k=0$. In fact, under these further conditions, we know from Assertion (v) that $\mathrm{E}(\infty,0)<\mathfrak{z}(0)=0$.
@@
By combining Equation (127) with the mean value theorem, we see that $\mathfrak{F}$ is Lipschitz equicontinuous. In particular this family of functions is equicontinuous. Moreover, it follows from (126) that $\mathfrak{F}$ is bounded in the supremum norm. We can hence apply the (Arzelà-) Ascoli theorem [44, Theorem A5], according to which $\partial_{x}\Phi(\mathrm{U}_{n},k_{0},\cdot)$, when restricted to the compact interval $[x_{0}-\delta,x_{0}+\delta]$, converges uniformly along some subsequence. Assume, for simplicity and without loss of generality, that the (full) sequence of functions converges itself. By Equation (118), recall that $\Phi(\mathrm{U},\cdot,\cdot)$ converges pointwise to the function $\Phi(\infty,\cdot,\cdot)$ on $\mathcal{O}$, which is is the real-valued function defined by (117). Then, by [93, Theorem 7.17],
@@
for every $(k,x)\in\mathcal{O}$. In the same way, we invoke Equations (124) and (125) together with the mean value theorem and the (Arzelà-) Ascoli theorem [44, Theorem A5] to deduce that
@@
This statement remains valid for $k=0$ provided that $\hat{\upsilon}(0)\neq 0$ and $r_{\mathfrak{p}}>0$ (i.e., $\mathfrak{p}_{1}\notin\mathbb{C}\mathfrak{e}_{0}$ or $\mathfrak{p}_{2}\notin\mathbb{C}\mathfrak{e}_{0}$).
@@
Note that $\mathrm{E}(\infty,k)\leq\mathfrak{b}\left(k\right)<\mathfrak{z}\left(k\right)$ for any $k\neq 0$. (When $\hat{\upsilon}(0)\neq 0$ and $\mathfrak{p}_{1}\notin\mathbb{C}\mathfrak{e}_{0}$ or $\mathfrak{p}_{2}\notin\mathbb{C}\mathfrak{e}_{0}$, we also have that $\mathrm{E}(\infty,0)<\mathfrak{b}(0)=\mathfrak{z}(0)$.) On the other hand, by Proposition 5.14, $\{(A_{1,1}(\mathrm{U},k)-\mathrm{E}(\infty,k)\mathfrak{1})^{-1}\}_{\mathrm{U}\geq 0}$ is a decreasing family of positive operators and, by Proposition 5.15, it converges strongly as $\mathrm{U}\rightarrow\infty$. Consequently, $(A_{1,1}(\mathrm{U},k)-\mathrm{E}(\mathrm{U},k)\mathfrak{1})^{-1}$ also converges strongly.
@@ 4.5 Spectral Gap and Anderson Localization
for any Hubbard coupling constant $\mathrm{U}\in\mathbb{R}_{0}^{+}$. When $r_{\mathfrak{p}}>0$ (i.e., $\mathfrak{p}_{1}\notin\mathbb{C}\mathfrak{e}_{0}$ or $\mathfrak{p}_{2}\notin\mathbb{C}\mathfrak{e}_{0}$) and $\hat{\upsilon}(0)\neq 0$, this quantity turns out to be strictly positive, uniformly with respect to the parameter $\mathrm{U}$:
@@
Fix $h_{b}\in[0,1/2]$. If $r_{\mathfrak{p}}>0$ (i.e., $\mathfrak{p}_{1}\notin\mathbb{C}\mathfrak{e}_{0}$ or $\mathfrak{p}_{2}\notin\mathbb{C}\mathfrak{e}_{0}$) and $\hat{\upsilon}(0)\neq 0$ then
@@
$\mathfrak{d}\left(k\right)\left(p\right)\doteq\mathfrak{\hat{p}}_{1}\left(k+p\right)+\mathfrak{\hat{p}}_{2}\left(k/2+p\right)\ ,\qquad p\in\mathbb{T}^{2}\ ,$ (136)
@@
Proof. By (36), we compute that
$\mathcal{F}^{-1}\left[(A_{1,1}(\mathrm{U},k)-\mathrm{E}(\mathrm{U},k)\mathfrak{1})^{-1}\mathfrak{d}\left(k\right)\right]$
$=(\mathcal{F}^{-1}A_{1,1}(\mathrm{U},k)\mathcal{F}-\mathrm{E}(\mathrm{U},k)\mathfrak{1})^{-1}\mathcal{F}^{-1}\left[\mathfrak{d}\left(k\right)\right]$
$={\sum_{y\in\mathbb{Z}^{2}}}\,\left(\mathrm{e}^{ik\cdot y}\mathfrak{p}_{1}(y)+\mathrm{e}^{i\frac{k}{2}\cdot y}\mathfrak{p}_{2}\left(y\right)\right)(\mathcal{F}^{-1}A_{1,1}(\mathrm{U},k)\mathcal{F}-\mathrm{E}(\mathrm{U},k)\mathfrak{1})^{-1}\mathfrak{e}_{y}\ .$ (138)
@@
$\left|\psi_{\mathrm{U},k}\left(x\right)\right|$ $=$ $\left|\left\langle\mathfrak{e}_{x},\psi_{\mathrm{U},k}\right\rangle\right|$ (143)
$\leq$ $\left|\hat{\upsilon}\left(k\right)\right|{\sum_{y\in\mathbb{Z}^{2}}}\left(\left|\mathfrak{p}_{1}\left(y\right)\right|+\left|\mathfrak{p}_{2}\left(y\right)\right|\right)\left|\left\langle\mathfrak{e}_{x},\left(\mathcal{F}^{-1}A_{1,1}\left(\mathrm{U},k\right)\mathcal{F}-\mathrm{E}(\mathrm{U},k)\mathfrak{1}\right)^{-1}\mathfrak{e}_{y}\right\rangle\right|$
$\leq$ $\frac{\left|\hat{\upsilon}\left(k\right)\right|}{\inf_{\mathrm{U}\in\mathbb{R}_{0}^{+}}\min_{k\in\mathbb{T}^{2}}\mathfrak{g}\left(\mathrm{U},k\right)-4\epsilon(\mathrm{e}^{\alpha}-1)}{\sum_{y\in\mathbb{Z}^{2}}}\left(\left|\mathfrak{p}_{1}\left(y\right)\right|+\left|\mathfrak{p}_{2}\left(y\right)\right|\right)\mathrm{e}^{-\alpha\left|x-y\right|}$
$\leq$ $\frac{\left|\hat{\upsilon}\left(k\right)\right|\mathrm{e}^{-\alpha\left|x\right|}}{\inf_{\mathrm{U}\in\mathbb{R}_{0}^{+}}\min_{k\in\mathbb{T}^{2}}\mathfrak{g}\left(\mathrm{U},k\right)-4\epsilon(\mathrm{e}^{\alpha}-1)}{\sum_{y\in\mathbb{Z}^{2}}}\left(\left|\mathfrak{p}_{1}\left(y\right)\right|+\left|\mathfrak{p}_{2}\left(y\right)\right|\right)\mathrm{e}^{\alpha\left|y\right|}$